OpenStudy (anonymous):

x+1/8=-3/4 What is x?

hartnn (hartnn):

-3/4-1/8

OpenStudy (anonymous):

-1/8 from both sides

OpenStudy (anonymous):

\[\frac{ x + 1 }{ 8 } = \frac{ -3 }{ 4 }\] (is this your equation?) multiplying by 2 on r.h.s \[\frac{ x + 1 }{ 8 } = \frac{ -6 }{ 8 }\] adding \[\frac{ -6 }{ 8 }\] \[\left( \frac{ x - 1 + 6 }{ 8 } \right) = 0\] x = -5